# Bregman Stochastic Proximal Point Algorithm with Variance Reduction

### v1

@@ 6 Experiments
We illustrate the advantage of variance reduction for BSAPA by performing in this section a numerical experiment on the Poisson linear inverse problem. It is given by:
@@
is the Kullback-Leibler divergence, $A\in\mathbb{R}^{n\times d}_{+}$ is the forward operator of the inverse problem, and $b\in\mathbb{R}_{++}^{n}$ is the measurements vector. It models the maximum likelihood estimation problem where the model is $b\sim\text{Poisson}(A{\mathsf{x}}_{*})$, with $\mathsf{x}$$_{*}$ the true unknown value. Each $f_{i}$ is convex. The kernel used is $h({\mathsf{x}})=-\sum_{i=0}^{d}\log\mathsf{x}_{i}$. Based on Bauschke et al. [3], we have that each function $f_{i}$ is $L$-relatively smooth w.r.t. $h$ when $L=\max_{i}b_{i}$. In these experiments, we compare a fixed stepsize BSAPA to non variance-reduced BSPPA with both fixed and vanishing stepsizes.
@@
using a gradient descent subroutine, since no closed-form solution is available in general. As this subroutine is employed in both algorithms under comparison, the evaluation remains equitable. Strictly speaking, this procedure corresponds to an approximate minimization of the proximal mapping, and hence the implementation is closer to an inexact variant of the algorithms. A formal treatment of inexactness lies beyond the scope of this work. Nevertheless, this approximation does not affect our empirical findings, which still clearly demonstrate the advantage of variance reduction. These findings encourage future extensions of the theoretical analysis toward inexact variants.
For the first experiment, we used a uniformly generated matrix. We set $n=500,d=100$. We couldn’t choose the sequence $(G_{k})_{k\in\mathbb{N}}$, as the existence of a gain function for the kernel is unknown. Following previous work [11] and their numerical experiments, we therefore assume $(G_{k})_{k\in\mathbb{N}}$ is constant, which can possibly go beyond our theoretical analysis. To account for this uncertainty and also the fact that $L$ is quite conservative, we consider several values for the stepsize in our experiments. Here, we present the results for two values; see Figure 1. For the complete range of stepsizes, see Appendix C, where all additional experimental results are provided. As expected, the results show that with vanishing stepsize (green curve), SPPA is slower that BSAPA (blue curve), and even plateaus quite early. With constant stepsize (orange curve), SPPA has the same rate as BSAPA (interpolation). However, while the stepsizes increase we observe that SPPA is less stable than BSAPA. More stability can be added to SPPA by using average iterates $(\bar{{x}})_{k\in\mathbb{N}}$. However this instability of SPPA, in the interpolation case, is vastly due to the approximate proximal mapping. Indeed when the $A$ is diagonal, a closed form exists for the proximal mapping and the experiments don’t exhibit this instability for SPPA; see Figure 2. Nevertheless, the experiments shows the advantage and stability of variance reduction, even in the interpolation case, when the proximal mapping can only be approximated.
Figure 1: Our BSAPA (variance reduced) is more stable, converges to the minimum and does not oscillate around it, even in the non-interpolation case with constant stepsize, contrary to BSPPA.
Figure 2: The case of diagonal AA: closed-form solution to the proximal mapping. There is no instability for BSPPA in that case, showing that its instability in Figure 1 is due to the inexact proximal mapping. This proves it is less stable to inexact algorithms compare to its variance-reduced counterparts.
@@
$\operatorname*{minimize}_{{\mathsf{x}}\in\mathbb{R}^{d}_{+}}\;\frac{1}{n}\sum_{i=1}^{n}\left\{f_{i}({\mathsf{x}})\coloneqq\mathsf{b}_{i}\log(\mathsf{b}_{i}/(A{\mathsf{x}})_{\theta_{i}})-\mathsf{b}_{i}+(A{\mathsf{x}})_{\theta_{i}}\right\}.$
The image used is the Shepp-Logan phantom. Also here, the vanishing stepsize SPPA is slower than BSAPA; see Figure 1. Furthermore, the results show, that constant stepsize SPPA oscillates around the minimum while BSAPA is stable; expected behaviors for non-interpolation cases. These oscillations are low for small stepsizes and high for bigger ones. We also ran this for different stepsizes because $G_{k}$ is unknown and $L$ is conservative.
Finally, in all experiments, BSAPA tends to explode quicker than SPPA when the stepsize is big enough. This is seen from the theory where SPPA does not have a bound on the stepsize; see Theorems 5.2 and 5.4. However bigger stepsizes increase the term which is related to the variance in those theorems, hence an increase in oscillations and no convergence. But until that stepsize threshold imposed by the theory, for non-interpolation cases, BSAPA remains more stable and ensure convergence whereas SPPA does not.
@@ 7 Conclusion
In this paper, we conducted a unified analysis of variance reduction for the Bregman stochastic proximal point algorithm (BSPPA). We proposed a generic variance-reduced algorithm based on BSPPA and prove convergence rates. From that general algorithm and analysis, we derived several new variance-reduced BSPPAs, employing SVRG- and SAGA-like techniques, and their corresponding convergence rates. More specifically, under the relative smoothness assumption, we prove sublinear and linear rates for convex and relatively strongly convex functions, respectively. Our general analysis can also recover the previously studied vanilla BSPPA with its standard rates. Furthermore, the unified theoretical results extend seamlessly to variance-reduced SGD with Bregman distance. For future work, we will focus on extending this work to nonsmooth functions and/or to inexact proximal mapping variants.
Acknowledgments.
P. Ochs acknowledges funding by the German Research Foundation for the project DFG Grant OC 150/3-1. The majority of this work was done while C. Traoré was a postdoc at Saarland University. He acknowledges the support of Occitanie region, the European Regional Development Fund (ERDF), and the French government, through the France 2030 project managed by the National Research Agency (ANR) with the reference number “ANR-22-EXES-0015”. C. Traoré thanks Edouard Pauwels for valuable discussions.
Appendices
@@ Appendix A Bregman SVRP (BSVRP)
Suppose that Assumptions (A.iii), 2.1 and (A.v) hold. Let $s\in\mathbb{N}$ and let $({x}_{k})_{k\in[m]}$ be the sequence generated by the inner iteration in Algorithm A.1. We assume that there exists a non-increasing sequence $(G_{k})_{k\in\mathbb{N}}$ such that, for all $i\in[n]$,
$G_{k}$ $\geq G\left(\nabla h({x}_{k}),\nabla h({x}_{k}),\frac{1}{L}(\nabla f_{i}({x}_{k})-\nabla f_{i}({\mathsf{x}}_{*}))\right),$ (A.1)
$G_{k}$ $\geq G\Big(\nabla h({x}_{k})-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}],\nabla h(\tilde{{x}}_{s}),\frac{1}{L}(\nabla f_{i}(\tilde{{x}}_{s})-\nabla f_{i}({\mathsf{x}}_{*}))\Big).$ (A.2)
Then, for all $k\in\{0,1,\cdots,m-1\}$,
$\mathsf{E}[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{s,k}]$ $\leq 2L\alpha_{k}^{2}G_{k}D_{F}({x}_{k},{\mathsf{x}}_{*})+\alpha_{k}^{2}G_{k}\sigma_{k}^{2}-\frac{1}{2}D_{h^{*}}\left(\nabla h({x}_{k}),\nabla h({x}_{k})+2\alpha_{k}\nabla F(\tilde{{x}}_{s})\right),$
@@
By Lemma A.2, Assumptions 4.5 are satisfied with $A_{k}=2LG_{k}$, $B_{k}=G_{k}$, $N_{k}=-(1/2\alpha_{k}^{2})D_{h^{*}}\left(\nabla h({x}_{k}),\nabla h({x}_{k})+2\alpha_{k}\nabla F(\tilde{{x}}_{s})\right)$, $\rho=C_{k}=0$. For simplicity, for any $k\in\mathbb{N}$, we set $A_{k}=A=2LG_{k}=2LG_{0}$, $B_{k}=B=G_{k}=G_{0}$, $\alpha_{k}=\alpha$.
Remark A.4.
For simplicity, for any $k\in\mathbb{N}$, we set $A_{k}=A=2LG_{k}=2LG_{0}$, $B_{k}=B=G_{k}=G_{0}$, $\alpha_{k}=\alpha$.
@@
We assume (A.iv) and that $\alpha<\frac{1}{4LG_{0}}$. Then, for all $s\in\mathbb{N}$ and under the assumptions of Lemma A.2,
@@ B.1 Fundamental results
Suppose that Assumptions 4.5 are verified, Assumption (A.ii) holds, and that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 4.1. Then, for all $k\in\mathbb{N}$,
@@
${x}_{k+1}$ $=\operatorname*{arg\,min}_{{\mathsf{x}}\in\mathbb{R}^{d}}\Big\{\underbrace{f_{i_{k}}({\mathsf{x}})-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle-\beta D_{h}({\mathsf{x}},{x}_{k})}_{:=R_{i_{k}}({\mathsf{x}})}+\left(\frac{1+\beta\alpha_{k}}{\alpha_{k}}\right)D_{h}({\mathsf{x}},{x}_{k})\Big\}$
@@
where ${{r}}_{k+1}\in\partial R({x}_{k+1})$ such that (B.1) holds.
@@
By using the three points identity in Lemma 2.5, $\langle\nabla h({x}_{k})-\nabla h({x}_{k+1}),{\mathsf{x}}-{x}_{k+1}\rangle=D_{h}({x}_{k+1},{x}_{k})+D_{h}({\mathsf{x}},{x}_{k+1})-D_{h}({\mathsf{x}},{x}_{k})$, in the previous inequality, we find
@@
Using the definitions of ${v}_{k}$ in (B.4) and (4.3), we can lower bound the left hand side as follows:
@@
$\qquad+f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}})$
$=\frac{1}{\alpha_{k}}\langle\nabla h({x}_{k})-\nabla h({z}_{k+1}),{x}_{k+1}-{x}_{k}\rangle$
@@
where in the first inequality, we used convexity of $f_{i}$, for all $i\in[n]$ and in the last one uses the three points inequality (2.5). By using (B.1) in (B.1) and $D_{h}({x}_{k+1},{z}_{k+1})\geq 0$, we obtain
@@
Now, define $\mathsf{E}_{k}[\cdot]=\mathsf{E}[\cdot\,|\,\mathfrak{F}_{k}]$, where $\mathfrak{F}_{k}$ is defined in Assumptions 4.5 and is such that ${x}_{k}$ is $\mathfrak{F}_{k}$-measurable and $i_{k}$ is independent of $\mathfrak{F}_{k}$. Thus, taking the conditional expectation in (B.1) and rearranging the terms, we have
@@
We show in the next proposition that the same result in Proposition B.1 stands for Bregman SGD with the generic variance reduction term ${e}_{k}$. As a consequence, since all the variance reduced results of the paper stem from Proposition B.1, they are also true for Bregman SGD with the same variance reduction techniques.
@@
Suppose that Assumptions 4.5 are verified, Assumption (A.ii) holds, and that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by the explicit version of Algorithm 4.1,i.e., SGD with the generic variance reduction term (4.4). Then, for all $k\in\mathbb{N}$,
$\left(1+\beta\alpha_{k}\right)\mathsf{E}[D_{h}({\mathsf{x}}_{*},{x}_{k+1})\,|\,\mathfrak{F}_{k}]$ $\coloneqq\left(1+\beta\alpha_{k}\right)\mathsf{E}[D_{h}({\mathsf{x}}_{*},{z}_{k+1})\,|\,\mathfrak{F}_{k}]$
@@
${x}_{k+1}$ $=\operatorname*{arg\,min}_{{\mathsf{x}}\in\mathbb{R}^{d}}\Big\{\underbrace{\langle{g}_{k},{\mathsf{x}}-{x}_{k}\rangle-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle-\beta D_{h}({\mathsf{x}},{x}_{k})}_{:=R_{i_{k}}({\mathsf{x}})}+\left(\frac{1+\beta\alpha_{k}}{\alpha_{k}}\right)D_{h}({\mathsf{x}},{x}_{k})\Big\}$
where ${g}_{k}\in\partial f_{i_{k}}({x}_{k})$ and ${{r}}_{k+1}\in\partial R({x}_{k+1})$ such that (B.7) holds.
Let ${\mathsf{x}}\in\mathcal{C}$. The function $R_{i_{k}}({\mathsf{x}})=\langle{g}_{k},{\mathsf{x}}-{x}_{k}\rangle-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle-\beta D_{h}({\mathsf{x}},{x}_{k})$ is convex, since $f_{i_{k}}$ is $\beta$-relatively strongly convex w.r.t. $h$. By convexity of $R_{i_{k}}$, we have:
@@
$=\langle{g}_{k},{\mathsf{x}}-{x}_{k}\rangle-\langle{g}_{k},{x}_{k+1}-{x}_{k}\rangle-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle+\langle{e}_{k},{x}_{k+1}-{x}_{k}\rangle$
$\qquad-\beta D_{h}({\mathsf{x}},{x}_{k})+\beta D_{h}({x}_{k+1},{x}_{k}).$ (B.8)
Since ${{r}}_{k+1}=\left(\frac{1+\beta\alpha_{k}}{\alpha_{k}}\right)\left(\nabla h({x}_{k})-\nabla h({x}_{k+1})\right)$ by (B.7), it follows from (B.1) and the convexity of $f_{i_{k}}$:
@@
$\qquad-\beta D_{h}({\mathsf{x}},{x}_{k})+\beta D_{h}({x}_{k+1},{x}_{k})$
$=f_{i_{k}}({\mathsf{x}})-f_{i_{k}}({x}_{k})-\langle{v}_{k},{x}_{k+1}-{x}_{k}\rangle-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle$
$\qquad-\beta D_{h}({\mathsf{x}},{x}_{k})+\beta D_{h}({x}_{k+1},{x}_{k}).$ (B.9)
By using the three points identity in Lemma 2.5, $\langle\nabla h({x}_{k})-\nabla h({x}_{k+1}),{\mathsf{x}}-{x}_{k+1}\rangle=D_{h}({x}_{k+1},{x}_{k})+D_{h}({\mathsf{x}},{x}_{k+1})-D_{h}({\mathsf{x}},{x}_{k})$, in the (B.1), we find
$\langle{v}_{k},{x}_{k+1}-{x}_{k}\rangle+f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}})+\frac{1}{\alpha_{k}}D_{h}({x}_{k+1},{x}_{k})$ $\leq\frac{1}{\alpha_{k}}D_{h}({\mathsf{x}},{x}_{k})-\frac{\left(1+\beta\alpha_{k}\right)}{\alpha_{k}}D_{h}({\mathsf{x}},{x}_{k+1})$
$\qquad-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle.$ (B.10)
@@
$=-\frac{1}{\alpha_{k}}D_{h}({x}_{k},{x}_{k+1})+f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}}),$ (B.12)
where in the second equality, we used convexity the three points inequality (2.5). By using (B.12) in (B.1), we obtain
$-\frac{1}{\alpha_{k}}D_{h}({x}_{k},{x}_{k+1})+f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}})$ $\leq\frac{1}{\alpha_{k}}D_{h}({\mathsf{x}},{x}_{k})-\frac{\left(1+\beta\alpha_{k}\right)}{\alpha_{k}}D_{h}({\mathsf{x}},{x}_{k+1})$
$\qquad-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle.$ (B.13)
Now, define $\mathsf{E}_{k}[\cdot]=\mathsf{E}[\cdot\,|\,\mathfrak{F}_{k}]$, where $\mathfrak{F}_{k}$ is defined in Assumptions 4.5 and is such that ${x}_{k}$ is $\mathfrak{F}_{k}$-measurable and $i_{k}$ is independent of $\mathfrak{F}_{k}$. Thus, taking the conditional expectation in (B.1) and rearranging the terms, we have
$\left(1+\beta\alpha_{k}\right)\mathsf{E}_{k}[D_{h}({\mathsf{x}},{x}_{k+1})]$ $\leq D_{h}({\mathsf{x}},{x}_{k})-\alpha_{k}\mathsf{E}_{k}\big[f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}})\big]+\mathsf{E}_{k}[D_{h}({x}_{k},{x}_{k+1})]$

### v2

@@ 6 Experiments
Figure 1: Our BSAPA (variance reduced) is more stable, converges to the minimum and does not oscillate around it, especially in the non-interpolation case with constant stepsize, contrary to BSPPA.
The following experiments go beyond the theory in certain aspects, which are pointed out. However, they still demonstrate the advantages of variance reduction. Using BSAPA, we illustrate those advantages by performing numerical experiments on the Poisson linear inverse problem given by:
@@
is the Kullback-Leibler divergence, $A\in\mathbb{R}^{n\times d}_{+}$ is the forward operator of the inverse problem, and $b\in\mathbb{R}_{++}^{n}$ is the measurements vector. It models the maximum likelihood estimation problem where the model is $b\sim\text{Poisson}(A{\mathsf{x}}_{*})$, with $\mathsf{x}$$_{*}$ the true unknown value. Each $f_{i}$ is convex. The kernel used is $h({\mathsf{x}})=-\sum_{i=0}^{d}\log\mathsf{x}_{i}$. Based on Bauschke et al. [2], we have that each function $f_{i}$ is $L$-relatively smooth w.r.t. $h$ when $L=\max_{i}b_{i}$. In addition, $\operatorname*{dom}h^{*}=(-\infty,0)$ which goes beyond the theoretical analysis. In these experiments, we compare a fixed stepsize BSAPA to non variance-reduced BSPPA with both fixed and vanishing stepsizes.
@@
using a gradient descent subroutine, since no closed-form solution is available in general, except when $A$ is diagonal. As this subroutine is employed in both algorithms under comparison, the evaluation remains equitable. Strictly speaking, this procedure corresponds to an approximate minimization of the proximal mapping, and hence the implementation is closer to an inexact variant of the algorithms. A formal treatment of inexactness lies beyond the scope of this work. Nevertheless, this approximation does not affect our empirical findings, which still clearly demonstrate the advantage of variance reduction. These findings encourage future extensions of the theoretical analysis toward inexact variants.
For the first experiment, we used a uniformly generated matrix. We set $n=500,d=100$. We couldn’t choose the sequence $(G_{k})_{k\in\mathbb{N}}$, as the gain function for the kernel is unknown exactly. Following previous work [11] and their numerical experiments, we therefore assume $(G_{k})_{k\in\mathbb{N}}$ is constant. To account for this uncertainty and also the fact that $L$ is quite conservative, we consider several values for the stepsize in our experiments. Here, we present the results for two values; see Figure 1(). For the complete range of stepsizes, see Appendix C(), where all additional experimental results are provided. As expected, the results show that with vanishing stepsize (green curve), SPPA is slower that BSAPA (blue curve), and even plateaus quite early. With constant stepsize (orange curve), SPPA has the same rate as BSAPA (interpolation). However, while the stepsizes increase we observe that SPPA is less stable than BSAPA. More stability can be added to SPPA by using average iterates $(\bar{{x}})_{k\in\mathbb{N}}$. This instability of SPPA, in the interpolation case, is vastly due to the approximate proximal mapping. Indeed when $A$ is diagonal, a closed form exists for the proximal mapping and the experiments don’t exhibit this instability for SPPA; see Figure 4(). Nevertheless, the experiments shows the advantage and stability of variance reduction, even in the interpolation case, when the proximal mapping can only be approximated.
@@
$\min_{{\mathsf{x}}\in\mathbb{R}^{d}_{+}}\frac{1}{n}\sum_{i=1}^{n}\Big\{f_{i}({\mathsf{x}})\coloneqq$ $\mathsf{b}_{i}\log(\mathsf{b}_{i}/(A{\mathsf{x}})_{\theta_{i}})-\mathsf{b}_{i}+(A{\mathsf{x}})_{\theta_{i}}\Big\}.$
The image used is the Shepp-Logan phantom. Also here, the vanishing stepsize SPPA is slower than BSAPA; see Figure 1(). Furthermore, the results show, that constant stepsize SPPA oscillates around the minimum while BSAPA is stable; expected behaviors for non-interpolation cases. These oscillations are low for small stepsizes and high for bigger ones. We also ran this for different stepsizes because $G_{k}$ is unknown and $L$ is conservative.
Finally, in all experiments, BSAPA tends to explode quicker than SPPA when the stepsize is big enough. This is seen from the theory where SPPA does not have a bound on the stepsize; see Theorems 5.2() and 5.4(). However bigger stepsizes increase the term which is related to the variance in those theorems, hence an increase in oscillations and no convergence. But until that stepsize threshold imposed by the theory, for non-interpolation cases, BSAPA remains more stable and ensure convergence whereas SPPA does not.
@@ 7 Conclusion
In this paper, we conducted a unified analysis of variance reduction for the Bregman stochastic proximal point algorithm (BSPPA). We proposed a generic variance-reduced algorithm based on BSPPA and prove convergence rates. From that general algorithm and analysis, we derived several new variance-reduced BSPPAs, employing SVRG- and SAGA-like techniques, and their corresponding convergence rates. More specifically, under the relative smoothness assumption, we prove sublinear and linear rates for convex and relatively strongly convex functions, respectively. Our general analysis can also recover the previously studied vanilla BSPPA with its standard rates. Furthermore, the unified theoretical results extend seamlessly to variance-reduced SGD with Bregman distance. Although not fully captured by the theory, the experiments demonstrated the advantages of variance reduction and highlighted the need for further refinement of the theoretical analysis. For future work, we will focus on extending this work to nonsmooth functions and/or to inexact proximal mapping variants.
Acknowledgements.
P. Ochs acknowledges funding by the German Research Foundation for the project DFG Grant OC 150/3-1. The majority of this work was done while C. Traoré was a postdoc at Saarland University. He acknowledges the support of Occitanie region, the European Regional Development Fund (ERDF), and the French government, through the France 2030 project managed by the National Research Agency (ANR) with the reference number “ANR-22-EXES-0015”.
@@ Appendix A Bregman SVRP (BSVRP)
Suppose that Assumptions (A.iii), 2.1() and (A.v) hold. We also assume that there exists a non-increasing sequence $(G_{\ell})_{\ell\in\mathbb{N}}$ such that, for all $i\in[n]$,
$G_{sm+k}$ $\geq G\left(\nabla h({x}_{k}),\nabla h({x}_{k}),\frac{1}{L}(\nabla f_{i}({x}_{k})-\nabla f_{i}({\mathsf{x}}_{*}))\right),$ (A.1)
$G_{sm+k}$ $\geq G\Big(\nabla h({x}_{k})-\mathsf{E}[\zeta_{k}\,|\,\mathfrak{F}_{s,k}],\nabla h(\tilde{{x}}_{s}),\frac{1}{L}(\nabla f_{i}(\tilde{{x}}_{s})-\nabla f_{i}({\mathsf{x}}_{*}))\Big),$ (A.2)
for any $s\in\mathbb{N}$, $k\in\{0,1,\cdots,m-1\}$ and $({x}_{k})_{k\in[m]}$ generated by the inner iteration of Algorithm A.1().
So let $s\in\mathbb{N}$ and let $({x}_{k})_{k\in[m]}$ be the sequence generated by the inner iteration in Algorithm A.1(). Then, for all $k\in\{0,1,\cdots,m-1\}$,
$\mathsf{E}[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{s,k}]$ $\leq 2L\alpha_{k}^{2}G_{sm+k}D_{F}({x}_{k},{\mathsf{x}}_{*})+\alpha_{k}^{2}G_{sm+k}\sigma_{k}^{2}-\frac{1}{2}D_{h^{*}}\left(\nabla h({x}_{k}),\nabla h({x}_{k})+2\alpha_{k}\nabla F(\tilde{{x}}_{s})\right),$
@@
By Lemma A.2(), Assumptions 4.5() are satisfied with $A_{k}=2LG_{sm+k}$, $B_{k}=G_{sm+k}$, $N_{k}=-(1/2\alpha_{k}^{2})D_{h^{*}}\left(\nabla h({x}_{k}),\nabla h({x}_{k})+2\alpha_{k}\nabla F(\tilde{{x}}_{s})\right)$, $\rho=C_{k}=0$. For simplicity, we set $A_{k}=A=2LG_{sm+k}=2LG_{0}$, $B_{k}=B=G_{sm+k}=G_{0}$, $\alpha_{k}=\alpha$.
@@
We assume (A.iv) and that $\alpha<\frac{1}{4LG_{0}}$. Then, for all $s\in\mathbb{N}$ and under the assumptions of Lemma A.2(),
@@ B.1 Fundamental results
Suppose that Assumptions 4.5() are verified, Assumption (A.ii) holds, and that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 4.1(). Then, for all $k\in\mathbb{N}$,
@@
${x}_{k+1}$ $=\operatorname*{arg\,min}_{{\mathsf{x}}\in\mathsf{H}}\Big\{\underbrace{f_{i_{k}}({\mathsf{x}})-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle-\beta D_{h}({\mathsf{x}},{x}_{k})}_{=:R_{i_{k}}({\mathsf{x}})}+\left(\frac{1+\beta\alpha_{k}}{\alpha_{k}}\right)D_{h}({\mathsf{x}},{x}_{k})\Big\}$
@@
where ${{r}}_{k+1}\in\partial R_{i_{k}}({x}_{k+1})$ such that (B.1) holds.
@@
By using the three points identity in Lemma 2.5(), $\langle\nabla h({x}_{k})-\nabla h({x}_{k+1}),{\mathsf{x}}-{x}_{k+1}\rangle=D_{h}({x}_{k+1},{x}_{k})+D_{h}({\mathsf{x}},{x}_{k+1})-D_{h}({\mathsf{x}},{x}_{k})$, in the previous inequality, we find
@@
Using (4.3), we can lower bound the left hand side as follows:
@@
$\qquad+f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}}),$
where in the first inequality, we used convexity of $f_{i}$, for all $i\in[n]$. Using the definitions of ${v}_{k}$ in (B.4), we can write
$-\langle{e}_{k},{x}_{k+1}-{x}_{k}\rangle+f_{i_{k}}({x}_{k+1})-f_{i_{k}}({\mathsf{x}})+\frac{1}{\alpha_{k}}D_{h}({x}_{k+1},{x}_{k})$ $=\frac{1}{\alpha_{k}}\langle\nabla h({x}_{k})-\nabla h({z}_{k+1}),{x}_{k+1}-{x}_{k}\rangle$
@@
where the last inequality uses the three points inequality (2.5). By using (B.1) in (B.1) and $D_{h}({x}_{k+1},{z}_{k+1})\geq 0$, we obtain
@@
Now, define $\mathsf{E}_{k}[\cdot]=\mathsf{E}[\cdot\,|\,\mathfrak{F}_{k}]$, where $\mathfrak{F}_{k}$ is defined in Assumptions 4.5() and is such that ${x}_{k}$ is $\mathfrak{F}_{k}$-measurable and $i_{k}$ is independent of $\mathfrak{F}_{k}$. Thus, taking the conditional expectation in (B.1) and rearranging the terms, we have
@@
We show in the next proposition that the same result in Proposition B.1() with $\beta=0$ stands for Bregman SGD with the generic variance reduction term ${e}_{k}$. As a consequence, since all the variance reduced results of the paper stem from Proposition B.1() with $\beta=0$, they are also true for Bregman SGD with the same variance reduction techniques.
@@
Suppose that Assumptions 4.5() are verified, and that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by the explicit version of Algorithm 4.1(),i.e., SGD with the generic variance reduction term (4.4). Then, for all $k\in\mathbb{N}$,
$\mathsf{E}[D_{h}({\mathsf{x}}_{*},{x}_{k+1})\,|\,\mathfrak{F}_{k}]$ $\coloneqq\mathsf{E}[D_{h}({\mathsf{x}}_{*},{z}_{k+1})\,|\,\mathfrak{F}_{k}]$
@@
${x}_{k+1}$ $=\operatorname*{arg\,min}_{{\mathsf{x}}\in\mathsf{H}}\Big\{\underbrace{\langle{g}_{k},{\mathsf{x}}-{x}_{k}\rangle-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle}_{=:R_{i_{k}}({\mathsf{x}})}+\frac{1}{\alpha_{k}}D_{h}({\mathsf{x}},{x}_{k})\Big\}$
$=\nabla h^{*}\left(\nabla h({x}_{k})-\alpha_{k}{{r}}_{k+1}\right),$ (B.7)
where ${g}_{k}\in\partial f_{i_{k}}({x}_{k})$ and ${{r}}_{k+1}=\nabla R_{i_{k}}({x}_{k+1})$ such that (B.7) holds.
Let ${\mathsf{x}}\in\mathcal{C}$. We have by convexity of $R_{i_{k}}$:
@@
$=\langle{g}_{k},{\mathsf{x}}-{x}_{k}\rangle-\langle{g}_{k},{x}_{k+1}-{x}_{k}\rangle$
$\qquad-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle+\langle{e}_{k},{x}_{k+1}-{x}_{k}\rangle.$ (B.8)
Since ${{r}}_{k+1}=\frac{1}{\alpha_{k}}\left(\nabla h({x}_{k})-\nabla h({x}_{k+1})\right)$ by (B.7), it follows from (B.1) and the convexity of $f_{i_{k}}$:
$\frac{1}{\alpha_{k}}\langle\nabla h({x}_{k})-\nabla h({x}_{k+1}),{\mathsf{x}}-{x}_{k+1}\rangle$
@@
$=f_{i_{k}}({\mathsf{x}})-f_{i_{k}}({x}_{k})-\langle{v}_{k},{x}_{k+1}-{x}_{k}\rangle-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle.$ (B.9)
By using the three points identity in Lemma 2.5(), $\langle\nabla h({x}_{k})-\nabla h({x}_{k+1}),{\mathsf{x}}-{x}_{k+1}\rangle=D_{h}({x}_{k+1},{x}_{k})+D_{h}({\mathsf{x}},{x}_{k+1})-D_{h}({\mathsf{x}},{x}_{k})$, in the (B.1), we find
$\langle{v}_{k},{x}_{k+1}-{x}_{k}\rangle+f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}})+\frac{1}{\alpha_{k}}D_{h}({x}_{k+1},{x}_{k})$
$\leq\frac{1}{\alpha_{k}}D_{h}({\mathsf{x}},{x}_{k})-\frac{1}{\alpha_{k}}D_{h}({\mathsf{x}},{x}_{k+1})-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle.$ (B.10)
@@
$=-\frac{1}{\alpha_{k}}D_{h}({x}_{k},{x}_{k+1})$
$\qquad+f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}}).$ (B.12)
By using (B.12) in (B.1), we obtain
$-\frac{1}{\alpha_{k}}D_{h}({x}_{k},{x}_{k+1})+f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}})$ $\leq\frac{1}{\alpha_{k}}D_{h}({\mathsf{x}},{x}_{k})-\frac{1}{\alpha_{k}}D_{h}({\mathsf{x}},{x}_{k+1})-\langle{e}_{k},{\mathsf{x}}-{x}_{k}\rangle.$ (B.13)
Now, define $\mathsf{E}_{k}[\cdot]=\mathsf{E}[\cdot\,|\,\mathfrak{F}_{k}]$, where $\mathfrak{F}_{k}$ is defined in Assumptions 4.5() and is such that ${x}_{k}$ is $\mathfrak{F}_{k}$-measurable and $i_{k}$ is independent of $\mathfrak{F}_{k}$. Thus, taking the conditional expectation in (B.13) and rearranging the terms, we have
$\mathsf{E}_{k}[D_{h}({\mathsf{x}},{x}_{k+1})]$ $\leq D_{h}({\mathsf{x}},{x}_{k})-\alpha_{k}\mathsf{E}_{k}\big[f_{i_{k}}({x}_{k})-f_{i_{k}}({\mathsf{x}})\big]+\mathsf{E}_{k}[D_{h}({x}_{k},{x}_{k+1})]$